OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (anonymous):

Do you know point slope form?

OpenStudy (anonymous):

use the formula y-y1=m(x-x1)

OpenStudy (anonymous):

Ya thats it^^

OpenStudy (anonymous):

y-(-1)=-1(x-4)